### v1

# Measurement of reactor antineutrino oscillation at SNO+ M. Abreu Laboratório de Instrumentação e Física Experimental de Partículas (LIP), Av. Prof. Gama Pinto, 2, 1649-003, Lisboa, Portugal Universidade de Lisboa, Instituto Superior Técnico (IST), Departamento de Física, Av. Rovisco Pais, 1049-001 Lisboa, Portugal V. Albanese

Measurement of reactor antineutrino oscillation at SNO+
M. Abreu
Laboratório de Instrumentação e Física Experimental de Partículas (LIP), Av. Prof. Gama Pinto, 2, 1649-003, Lisboa, Portugal Universidade de Lisboa, Instituto Superior Técnico (IST), Departamento de Física, Av. Rovisco Pais, 1049-001 Lisboa, Portugal V. Albanese
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada A. Allega
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada R. Alves
Laboratório de Instrumentação e Física Experimental de Partículas (LIP), 3004-516, Coimbra, Portugal M. R. Anderson
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada S. Andringa
Laboratório de Instrumentação e Física Experimental de Partículas (LIP), Av. Prof. Gama Pinto, 2, 1649-003, Lisboa, Portugal L. Anselmo
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada J. Antunes
Laboratório de Instrumentação e Física Experimental de Partículas (LIP), Av. Prof. Gama Pinto, 2, 1649-003, Lisboa, Portugal Universidade de Lisboa, Instituto Superior Técnico (IST), Departamento de Física, Av. Rovisco Pais, 1049-001 Lisboa, Portugal E. Arushanova
Queen Mary, University of London, School of Physics and Astronomy, 327 Mile End Road, London, E1 4NS, UK S. Asahi
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada M. Askins
University of California, Berkeley, Department of Physics, CA 94720, Berkeley, USA Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720-8153, USA University of California, Davis, 1 Shields Avenue, Davis, CA 95616, USA D. M. Asner
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada Brookhaven National Laboratory, P.O. Box 5000, Upton, NY 11973-500, USA D. J. Auty
University of Alberta, Department of Physics, 4-181 CCIS, Edmonton, AB T6G 2E1, Canada A. R. Back
Queen Mary, University of London, School of Physics and Astronomy, 327 Mile End Road, London, E1 4NS, UK University of Sussex, Physics & Astronomy, Pevensey II, Falmer, Brighton, BN1 9QH, UK S. Back
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada A. Bacon
University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA T. Baltazar
Laboratório de Instrumentação e Física Experimental de Partículas (LIP), Av. Prof. Gama Pinto, 2, 1649-003, Lisboa, Portugal Universidade de Lisboa, Instituto Superior Técnico (IST), Departamento de Física, Av. Rovisco Pais, 1049-001 Lisboa, Portugal F. Barão
Laboratório de Instrumentação e Física Experimental de Partículas (LIP), Av. Prof. Gama Pinto, 2, 1649-003, Lisboa, Portugal Universidade de Lisboa, Instituto Superior Técnico (IST), Departamento de Física, Av. Rovisco Pais, 1049-001 Lisboa, Portugal Z. Barnard
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada A. Barr
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada N. Barros
Laboratório de Instrumentação e Física Experimental de Partículas, Rua Larga, 3004-516 Coimbra, Portugal Universidade de Coimbra, Departamento de Física (FCTUC), 3004-516, Coimbra, Portugal University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA Technische Universität Dresden, Institut für Kern und Teilchenphysik, Zellescher Weg 19, Dresden, 01069, Germany Laboratório de Instrumentação e Física Experimental de Partículas (LIP), Av. Prof. Gama Pinto, 2, 1649-003, Lisboa, Portugal D. Bartlett
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada R. Bayes
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada C. Beaudoin
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada E. W. Beier
University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA G. Berardi
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada T. S. Bezerra
University of Sussex, Physics & Astronomy, Pevensey II, Falmer, Brighton, BN1 9QH, UK A. Bialek
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada University of Alberta, Department of Physics, 4-181 CCIS, Edmonton, AB T6G 2E1, Canada S. D. Biller
University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK E. Blucher
The Enrico Fermi Institute and Department of Physics, The University of Chicago, Chicago, IL 60637, USA A. Boeltzig
Technische Universität Dresden, Institut für Kern und Teilchenphysik, Zellescher Weg 19, Dresden, 01069, Germany R. Bonventre
University of California, Berkeley, Department of Physics, CA 94720, Berkeley, USA Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720-8153, USA University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA M. Boulay
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada D. Braid
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada E. Caden
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada E. J. Callaghan
University of California, Berkeley, Department of Physics, CA 94720, Berkeley, USA Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720-8153, USA University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA J. Caravaca
University of California, Berkeley, Department of Physics, CA 94720, Berkeley, USA Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720-8153, USA J. Carvalho
Laboratório de Instrumentação e Física Experimental de Partículas (LIP), 3004-516, Coimbra, Portugal Universidade de Coimbra, Departamento de Física (FCTUC), 3004-516, Coimbra, Portugal L. Cavalli
University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK D. Chauhan
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada Laboratório de Instrumentação e Física Experimental de Partículas (LIP), Av. Prof. Gama Pinto, 2, 1649-003, Lisboa, Portugal Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada M. Chen
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada S. Cheng
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada O. Chkvorets
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada K. J. Clark
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada University of Sussex, Physics & Astronomy, Pevensey II, Falmer, Brighton, BN1 9QH, UK University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK B. Cleveland
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada C. Connors
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada D. Cookman
King’s College London, Department of Physics, Strand Building, Strand, London, WC2R 2LS, UK University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK J. Corning
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada I. T. Coulter
University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK M. A. Cox
University of Liverpool, Department of Physics, Liverpool, L69 3BX, UK Laboratório de Instrumentação e Física Experimental de Partículas (LIP), Av. Prof. Gama Pinto, 2, 1649-003, Lisboa, Portugal D. Cressy
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada X. Dai
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada C. Darrach
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada S. DeGraw
University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK R. Dehghani
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada J. Deloye
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada M. M. Depatie
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada F. Descamps
University of California, Berkeley, Department of Physics, CA 94720, Berkeley, USA Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720-8153, USA C. Dima
University of Sussex, Physics & Astronomy, Pevensey II, Falmer, Brighton, BN1 9QH, UK J. Dittmer
Technische Universität Dresden, Institut für Kern und Teilchenphysik, Zellescher Weg 19, Dresden, 01069, Germany K. H. Dixon
King’s College London, Department of Physics, Strand Building, Strand, London, WC2R 2LS, UK F. Di Lodovico
King’s College London, Department of Physics, Strand Building, Strand, London, WC2R 2LS, UK Queen Mary, University of London, School of Physics and Astronomy, 327 Mile End Road, London, E1 4NS, UK A. Doxtator
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada N. Duhaime
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada F. Duncan
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada J. Dunger
Queen Mary, University of London, School of Physics and Astronomy, 327 Mile End Road, London, E1 4NS, UK University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK A. D. Earle
University of Sussex, Physics & Astronomy, Pevensey II, Falmer, Brighton, BN1 9QH, UK M. S. Esmaeilian
University of Alberta, Department of Physics, 4-181 CCIS, Edmonton, AB T6G 2E1, Canada D. Fabris
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada E. Falk
University of Sussex, Physics & Astronomy, Pevensey II, Falmer, Brighton, BN1 9QH, UK A. Farrugia
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada N. Fatemighomi
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada C. Felber
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada V. Fischer
University of California, Davis, 1 Shields Avenue, Davis, CA 95616, USA E. Fletcher
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada R. Ford
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada K. Frankiewicz
Boston University, Department of Physics, 590 Commonwealth Avenue, Boston, MA 02215, USA N. Gagnon
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada A. Gaur
University of Alberta, Department of Physics, 4-181 CCIS, Edmonton, AB T6G 2E1, Canada J. Gauthier
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada A. Gibson-Foster
University of Sussex, Physics & Astronomy, Pevensey II, Falmer, Brighton, BN1 9QH, UK K. Gilje
University of Alberta, Department of Physics, 4-181 CCIS, Edmonton, AB T6G 2E1, Canada O. I. González-Reina
Universidad Nacional Autónoma de México (UNAM), Instituto de Física, Apartado Postal 20-364, México D.F., 01000, México D. Gooding
Boston University, Department of Physics, 590 Commonwealth Avenue, Boston, MA 02215, USA P. Gorel
University of Alberta, Department of Physics, 4-181 CCIS, Edmonton, AB T6G 2E1, Canada K. Graham
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada C. Grant
Boston University, Department of Physics, 590 Commonwealth Avenue, Boston, MA 02215, USA University of California, Davis, 1 Shields Avenue, Davis, CA 95616, USA J. Grove
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada S. Grullon
University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA E. Guillian
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada R. L. Hahn
Brookhaven National Laboratory, P.O. Box 5000, Upton, NY 11973-500, USA S. Hall
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada A. L. Hallin
University of Alberta, Department of Physics, 4-181 CCIS, Edmonton, AB T6G 2E1, Canada D. Hallman
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada S. Hans
Brookhaven National Laboratory, P.O. Box 5000, Upton, NY 11973-500, USA J. Hartnell
University of Sussex, Physics & Astronomy, Pevensey II, Falmer, Brighton, BN1 9QH, UK P. Harvey
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada C. Hearns
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada M. R. Hebert
University of California, Berkeley, Department of Physics, CA 94720, Berkeley, USA Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720-8153, USA M. Hedayatipour
University of Alberta, Department of Physics, 4-181 CCIS, Edmonton, AB T6G 2E1, Canada W. J. Heintzelman
University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA J. Heise
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada R. L. Helmer
TRIUMF, 4004 Wesbrook Mall, Vancouver, BC V6T 2A3, Canada C. Hewitt
University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK B. Hodak
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada M. Hodak
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada M. Hood
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada D. Horne
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada M. Howe
University of North Carolina, Department of Physics and Astronomy, Phillips Hall, Chapell Hill, NC 27599-3255, USA B. Hreljac
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada J. Hu
University of Alberta, Department of Physics, 4-181 CCIS, Edmonton, AB T6G 2E1, Canada P. Huang
University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK R. Hunt-Stokes
University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK T. Iida
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada A. S. Inácio
University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK Laboratório de Instrumentação e Física Experimental de Partículas (LIP), Av. Prof. Gama Pinto, 2, 1649-003, Lisboa, Portugal Universidade de Lisboa, Faculdade de Ciências (FCUL), Departamento de Física, Campo Grande, Edifício C8, 1749-016 Lisboa, Portugal C. M. Jackson
University of California, Berkeley, Department of Physics, CA 94720, Berkeley, USA Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720-8153, USA N. A. Jelley
University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK C. J. Jillings
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada C. Jones
University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK P. G. Jones
Queen Mary, University of London, School of Physics and Astronomy, 327 Mile End Road, London, E1 4NS, UK University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK S. Kaluzienski
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada K. Kamdin
University of California, Berkeley, Department of Physics, CA 94720, Berkeley, USA Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720-8153, USA T. Kaptanoglu
University of California, Berkeley, Department of Physics, CA 94720, Berkeley, USA Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720-8153, USA University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA J. Kaspar
University of Washington, Center for Experimental Nuclear Physics and Astrophysics, and Department of Physics, Seattle, WA 98195, USA K. Keeter
Idaho State University, 921 S. 8th Ave, Mail Stop 8106, Pocatello, ID 83209-8106 C. Kefelian
University of California, Berkeley, Department of Physics, CA 94720, Berkeley, USA Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720-8153, USA P. Khaghani
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada L. Kippenbrock
University of Washington, Center for Experimental Nuclear Physics and Astrophysics, and Department of Physics, Seattle, WA 98195, USA J. Kladnik
Laboratório de Instrumentação e Física Experimental de Partículas (LIP), Av. Prof. Gama Pinto, 2, 1649-003, Lisboa, Portugal J. R. Klein
University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA R. Knapik
University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA Norwich University, 158 Harmon Drive, Northfield, VT 05663, USA J. Kofron
University of Washington, Center for Experimental Nuclear Physics and Astrophysics, and Department of Physics, Seattle, WA 98195, USA L. L. Kormos
Lancaster University, Physics Department, Lancaster, LA1 4YB, UK S. Korte
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada B. Krar
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada C. Kraus
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada C. B. Krauss
University of Alberta, Department of Physics, 4-181 CCIS, Edmonton, AB T6G 2E1, Canada T. Kroupová
University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK K. Labe
The Enrico Fermi Institute and Department of Physics, The University of Chicago, Chicago, IL 60637, USA F. Lafleur
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada C. Lake
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada I. Lam
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada C. Lan
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada B. J. Land
University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA University of California, Berkeley, Department of Physics, CA 94720, Berkeley, USA Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720-8153, USA R. Lane
Queen Mary, University of London, School of Physics and Astronomy, 327 Mile End Road, London, E1 4NS, UK S. Langrock
Queen Mary, University of London, School of Physics and Astronomy, 327 Mile End Road, London, E1 4NS, UK P. Larochelle
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada S. Larose
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada A. LaTorre
The Enrico Fermi Institute and Department of Physics, The University of Chicago, Chicago, IL 60637, USA I. Lawson
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada L. Lebanowski
University of California, Berkeley, Department of Physics, CA 94720, Berkeley, USA Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720-8153, USA University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA J. Lee
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada C. Lefebvre
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada G. M. Lefeuvre
University of Sussex, Physics & Astronomy, Pevensey II, Falmer, Brighton, BN1 9QH, UK E. J. Leming
University of Sussex, Physics & Astronomy, Pevensey II, Falmer, Brighton, BN1 9QH, UK University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK A. Li
Boston University, Department of Physics, 590 Commonwealth Avenue, Boston, MA 02215, USA O. Li
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada J. Lidgard
University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK B. Liggins
Queen Mary, University of London, School of Physics and Astronomy, 327 Mile End Road, London, E1 4NS, UK P. Liimatainen
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada Y. H. Lin
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada X. Liu
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada Y. Liu
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada V. Lozza
Laboratório de Instrumentação e Física Experimental de Partículas (LIP), Av. Prof. Gama Pinto, 2, 1649-003, Lisboa, Portugal Universidade de Lisboa, Faculdade de Ciências (FCUL), Departamento de Física, Campo Grande, Edifício C8, 1749-016 Lisboa, Portugal Technische Universität Dresden, Institut für Kern und Teilchenphysik, Zellescher Weg 19, Dresden, 01069, Germany M. Luo
University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA S. Maguire
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada Brookhaven National Laboratory, P.O. Box 5000, Upton, NY 11973-500, USA Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada A. Maio
Laboratório de Instrumentação e Física Experimental de Partículas (LIP), Av. Prof. Gama Pinto, 2, 1649-003, Lisboa, Portugal Universidade de Lisboa, Faculdade de Ciências (FCUL), Departamento de Física, Campo Grande, Edifício C8, 1749-016 Lisboa, Portugal K. Majumdar
University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK S. Manecki
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada J. Maneira
Laboratório de Instrumentação e Física Experimental de Partículas (LIP), Av. Prof. Gama Pinto, 2, 1649-003, Lisboa, Portugal Universidade de Lisboa, Faculdade de Ciências (FCUL), Departamento de Física, Campo Grande, Edifício C8, 1749-016 Lisboa, Portugal R. D. Martin
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada E. Marzec
University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA A. Mastbaum
The Enrico Fermi Institute and Department of Physics, The University of Chicago, Chicago, IL 60637, USA University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA A. Mathewson
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada N. McCauley
University of Liverpool, Department of Physics, Liverpool, L69 3BX, UK A. B. McDonald
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada K. McFarlane
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada P. Mekarski
University of Alberta, Department of Physics, 4-181 CCIS, Edmonton, AB T6G 2E1, Canada M. Meyer
Technische Universität Dresden, Institut für Kern und Teilchenphysik, Zellescher Weg 19, Dresden, 01069, Germany C. Miller
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada C. Mills
University of Sussex, Physics & Astronomy, Pevensey II, Falmer, Brighton, BN1 9QH, UK G. Milton
University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK M. Mlejnek
University of Sussex, Physics & Astronomy, Pevensey II, Falmer, Brighton, BN1 9QH, UK E. Mony
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada B. Morissette
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada D. Morris
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada I. Morton-Blake
University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK M. J. Mottram
University of Sussex, Physics & Astronomy, Pevensey II, Falmer, Brighton, BN1 9QH, UK Queen Mary, University of London, School of Physics and Astronomy, 327 Mile End Road, London, E1 4NS, UK M. Mubasher
University of Alberta, Department of Physics, 4-181 CCIS, Edmonton, AB T6G 2E1, Canada S. Nae
Laboratório de Instrumentação e Física Experimental de Partículas (LIP), Av. Prof. Gama Pinto, 2, 1649-003, Lisboa, Portugal Universidade de Lisboa, Faculdade de Ciências (FCUL), Departamento de Física, Campo Grande, Edifício C8, 1749-016 Lisboa, Portugal S. Naugle
University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA M. Newcomer
University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA M. Nirkko
University of Sussex, Physics & Astronomy, Pevensey II, Falmer, Brighton, BN1 9QH, UK L. J. Nolan
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada Queen Mary, University of London, School of Physics and Astronomy, 327 Mile End Road, London, E1 4NS, UK V. M. Novikov
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada H. M. O’Keeffe
Lancaster University, Physics Department, Lancaster, LA1 4YB, UK Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada E. O’Sullivan
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada G. D. Orebi Gann
University of California, Berkeley, Department of Physics, CA 94720, Berkeley, USA Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720-8153, USA University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA S. Ouyang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China J. Page
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada University of Sussex, Physics & Astronomy, Pevensey II, Falmer, Brighton, BN1 9QH, UK S. Pal
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada K. Paleshi
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada W. Parker
University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK M. J. Parnell
Lancaster University, Physics Department, Lancaster, LA1 4YB, UK J. Paton
University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK S. J. M. Peeters
University of Sussex, Physics & Astronomy, Pevensey II, Falmer, Brighton, BN1 9QH, UK T. Pershing
University of California, Davis, 1 Shields Avenue, Davis, CA 95616, USA Z. Petriw
University of Alberta, Department of Physics, 4-181 CCIS, Edmonton, AB T6G 2E1, Canada J. Petzoldt
Technische Universität Dresden, Institut für Kern und Teilchenphysik, Zellescher Weg 19, Dresden, 01069, Germany L. J. Pickard
University of California, Berkeley, Department of Physics, CA 94720, Berkeley, USA Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720-8153, USA University of California, Davis, 1 Shields Avenue, Davis, CA 95616, USA D. Pracsovics
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada G. Prior
Laboratório de Instrumentação e Física Experimental de Partículas (LIP), Av. Prof. Gama Pinto, 2, 1649-003, Lisboa, Portugal J. C. Prouty
University of California, Berkeley, Department of Physics, CA 94720, Berkeley, USA Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720-8153, USA B. Quenallata
Laboratório de Instrumentação e Física Experimental de Partículas, Rua Larga, 3004-516 Coimbra, Portugal Universidade de Coimbra, Departamento de Física (FCTUC), 3004-516, Coimbra, Portugal S. Quirk
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada P. Ravi
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada S. Read
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada A. Reichold
University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK M. Reinhard
Technische Universität Dresden, Institut für Kern und Teilchenphysik, Zellescher Weg 19, Dresden, 01069, Germany S. Riccetto
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada M. Rigan
University of Sussex, Physics & Astronomy, Pevensey II, Falmer, Brighton, BN1 9QH, UK I. Ritchie
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada A. Robertson
University of Liverpool, Department of Physics, Liverpool, L69 3BX, UK B. C. Robertson
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada J. Rose
University of Liverpool, Department of Physics, Liverpool, L69 3BX, UK R. Rosero
Brookhaven National Laboratory, P.O. Box 5000, Upton, NY 11973-500, USA P. M. Rost
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada J. Rumleskie
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada A. Sörensen
Technische Universität Dresden, Institut für Kern und Teilchenphysik, Zellescher Weg 19, Dresden, 01069, Germany P. Schrock
Technische Universität Dresden, Institut für Kern und Teilchenphysik, Zellescher Weg 19, Dresden, 01069, Germany M. A. Schumaker
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada M. H. Schwendener
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada D. Scislowski
University of Washington, Center for Experimental Nuclear Physics and Astrophysics, and Department of Physics, Seattle, WA 98195, USA J. Secrest
University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA Armstrong Atlantic State University, 11935 Abercorn Street, Savannah, GA 31419, USA M. Seddighin
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada L. Segui
University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK S. Seibert
University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA I. Semenec
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada F. Shaker
University of Alberta, Department of Physics, 4-181 CCIS, Edmonton, AB T6G 2E1, Canada T. Shantz
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada M. K. Sharma
University of Alberta, Department of Chemistry, 1-001 CCIS, Edmonton, AB T6G 2E9, Canada J. Shen
University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA T. M. Shokair
University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA L. Sibley
University of Alberta, Department of Physics, 4-181 CCIS, Edmonton, AB T6G 2E1, Canada J. Simms
University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK J. R. Sinclair
University of Sussex, Physics & Astronomy, Pevensey II, Falmer, Brighton, BN1 9QH, UK K. Singh
University of Alberta, Department of Physics, 4-181 CCIS, Edmonton, AB T6G 2E1, Canada P. Skensved
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada M. Smiley
University of California, Berkeley, Department of Physics, CA 94720, Berkeley, USA Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720-8153, USA T. Sonley
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada M. St-Amant
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada R. Stainforth
University of Liverpool, Department of Physics, Liverpool, L69 3BX, UK S. Stankiewicz
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada M. Strait
The Enrico Fermi Institute and Department of Physics, The University of Chicago, Chicago, IL 60637, USA M. I. Stringer
Queen Mary, University of London, School of Physics and Astronomy, 327 Mile End Road, London, E1 4NS, UK University of Sussex, Physics & Astronomy, Pevensey II, Falmer, Brighton, BN1 9QH, UK A. Stripay
SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada R. Svoboda
University of California, Davis, 1 Shields Avenue, Davis, CA 95616, USA S. Tacchino
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada R. Tafirout
TRIUMF, 4004 Wesbrook Mall, Vancouver, BC V6T 2A3, Canada B. Tam
University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada C. Tanguay
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada J. Tatar
University of Washington, Center for Experimental Nuclear Physics and Astrophysics, and Department of Physics, Seattle, WA 98195, USA L. Tian
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada N. Tolich
University of Washington, Center for Experimental Nuclear Physics and Astrophysics, and Department of Physics, Seattle, WA 98195, USA J. Tseng
University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK H. W. C. Tseung
University of Washington, Center for Experimental Nuclear Physics and Astrophysics, and Department of Physics, Seattle, WA 98195, USA E. Turner
University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK E. Vázquez-Jáuregui
Universidad Nacional Autónoma de México (UNAM), Instituto de Física, Apartado Postal 20-364, México D.F., 01000, México SNOLAB, Creighton Mine #9, 1039 Regional Road 24, Sudbury, ON P3Y 1N2, Canada Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada S. Valder
University of Sussex, Physics & Astronomy, Pevensey II, Falmer, Brighton, BN1 9QH, UK R. Van Berg
University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA J. G. C. Veinot
University of Alberta, Department of Chemistry, 11227 Saskatchewan Drive, Edmonton, Alberta, T6G 2G2, Canada C. J. Virtue
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada B. von Krosigk
Technische Universität Dresden, Institut für Kern und Teilchenphysik, Zellescher Weg 19, Dresden, 01069, Germany J. M. G. Walker
University of Liverpool, Department of Physics, Liverpool, L69 3BX, UK M. Walker
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada J. Wallig
Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720-8153, USA S. C. Walton
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada F. Wang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China J. Wang
University of Oxford, The Denys Wilkinson Building, Keble Road, Oxford, OX1 3RH, UK M. Ward
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada J. Waterfield
University of Sussex, Physics & Astronomy, Pevensey II, Falmer, Brighton, BN1 9QH, UK J. J. Weigand
Technische Universität Dresden, Faculty of Chemistry and Food Chemistry, Dresden, 01062, Germany R. F. White
University of Sussex, Physics & Astronomy, Pevensey II, Falmer, Brighton, BN1 9QH, UK J. F. Wilkerson
University of North Carolina, Department of Physics and Astronomy, Phillips Hall, Chapell Hill, NC 27599-3255, USA J. R. Wilson
King’s College London, Department of Physics, Strand Building, Strand, London, WC2R 2LS, UK Queen Mary, University of London, School of Physics and Astronomy, 327 Mile End Road, London, E1 4NS, UK J. D. Wilson
University of Alberta, Department of Physics, 4-181 CCIS, Edmonton, AB T6G 2E1, Canada T. J. Winchester
University of Washington, Center for Experimental Nuclear Physics and Astrophysics, and Department of Physics, Seattle, WA 98195, USA P. Woosaree
Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada A. Wright
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada S. Yang
University of Alberta, Department of Physics, 4-181 CCIS, Edmonton, AB T6G 2E1, Canada K. Yazigi
University of Alberta, Department of Physics, 4-181 CCIS, Edmonton, AB T6G 2E1, Canada Z. Ye
University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA M. Yeh
Brookhaven National Laboratory, P.O. Box 5000, Upton, NY 11973-500, USA S. Yu
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada Laurentian University, School of Natural Sciences, 935 Ramsey Lake Road, Sudbury, ON P3E 2C6, Canada T. Zhang
University of California, Davis, 1 Shields Avenue, Davis, CA 95616, USA Y. Zhang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China University of Alberta, Department of Physics, 4-181 CCIS, Edmonton, AB T6G 2E1, Canada T. Zhao
Queen’s University, Department of Physics, Engineering Physics & Astronomy, Kingston, ON K7L 3N6, Canada K. Zuber
Technische Universität Dresden, Institut für Kern und Teilchenphysik, Zellescher Weg 19, Dresden, 01069, Germany MTA Atomki, 4001 Debrecen, Hungary A. Zummo
University of Pennsylvania, Department of Physics & Astronomy, 209 South 33rd Street, Philadelphia, PA 19104-6396, USA
(May 7, 2025)
Abstract
The SNO+ collaboration reports its second spectral analysis of reactor antineutrino oscillation using 286 tonne-years of new data. The measured energies of reactor antineutrino candidates were fitted to obtain the second-most precise determination of the neutrino mass-squared difference = () 10-5 eV2.
Constraining and with measurements from long-baseline reactor antineutrino and solar neutrino experiments yields = () 10-5 eV2 and . This fit also yields a first measurement of the flux of geoneutrinos in the Western Hemisphere, with TNU at SNO+.
Introduction.
Neutrino oscillation is well established through measurements of neutrino rates and energy spectra from particle accelerators, nuclear reactors, the atmosphere, and the Sun. These measurements provide consistent values for the three neutrino oscillation angles and two mass-squared differences , where and are 1, 2, or 3 ().
Leading measurements of , , and have been obtained with reactor antineutrinos [1, 2, 3, 4].
The KamLAND experiment has made the most precise measurement of , along with a less sensitive measurement of [1]. In contrast, solar neutrino experiments have provided the most precise measurement of while being less sensitive to . The KamLAND measurement of is currently in 1.5 tension with the result of a combined analysis of all available solar neutrino data performed by Super-K [5].
SNO+ has previously published results studying reactor antineutrinos, including the first evidence of reactor antineutrino detection with a water Cherenkov detector [6] and an initial measurement of reactor antineutrino oscillation with the detector nearly half-filled with liquid scintillator [7].
Radioactive decays within the Earth also produce antineutrinos, which have been observed by the KamLAND and Borexino liquid scintillator detectors in Japan and Italy, respectively [8, 9].
This Letter presents a measurement of and using reactor antineutrinos, and initial constraints on the geoneutrino flux, based on the first data collected by SNO+ as a fully-filled liquid scintillator detector.
Data.
The SNO+ detector now contains about 780 tonnes of liquid scintillator within its 6.0-m radius spherical acrylic vessel (AV). Light produced by interactions in the scintillator is detected by 9362 inward-facing photomultiplier tubes (PMTs) at a radius of about 8.5 m. The scintillator volume is shielded by ultrapure water between the AV and PMTs and also beyond, where outward-looking PMTs are used to detect muons.
The SNO+ detector is described in detail in Ref. [10].
The livetime of the dataset used in this analysis is 134.4 days, collected between May 17, 2022, and March 14, 2023. During this period, the scintillator was linear alkylbenzene (LAB) with 2,5-Diphenyloxazole (PPO) at a concentration of 2.2 g/L.
The collected scintillation light yield was measured to be about 210 ‘clean’ PMT hits/MeV, which depends on the number of working channels and includes the removal of hits due to electronic noise and exclusion of channels that are not considered well calibrated in terms of charge and timing. The trigger threshold for this dataset was around 20 PMT hits, corresponding to approximately 0.09 MeV, well below the energy of the reactor antineutrino signals of interest.
The properties of particle interactions are inferred using the times and locations of hit PMTs.
The time of flight of the photons detected by the PMTs is used to reconstruct the interaction position.
The number of hit PMTs is approximately proportional to the energy deposited by a particle. The position-dependent efficiency to detect photons and the probability of detecting multiple photons on individual PMTs are accounted for in the energy reconstruction using Monte Carlo (MC) simulations.
The resolution of the reconstructed energy is about 6.5%/ for an electron at the center of the detector in this dataset. The reconstructed position resolution for a 2.5-MeV electron at the center is 12 cm in each of the three Cartesian axes.
Calibrations.
Calibrations were performed using 214Bi decays and 214Po decays identified by their delayed coincidence. These decays were sourced by the 238U intrinsic to the detector as well as ingress of 222Rn into the scintillator. The selection criteria described in Table 1, which include the time () and distance () between reconstructed interactions, provide a highly-pure sample of 214BiPo coincidences.
Table 1: Selection criteria applied to data and simulations to select 214BiPo and Reactor- IBD coincidence events. See text for parameter definitions.
214BiPo
Reactor- IBD
Prompt
Delayed
Prompt
Delayed
1.25-3.0
0.7-1.1
0.9-9.0
1.85-2.5
0-4.0
0-4.0
0-5.7
0-5.7
0-1.0
0-2.5
3.7-1000
0-2000
LR
-3.5
The time profile of the scintillator is modeled for both and particles as a sum of exponentials with a single rise time parameter. The decay constants and amplitudes are tuned to provide the best match between the time-of-flight-corrected hit times in data and Monte Carlo (MC) simulation. This calibration is performed for ’s and ’s separately since ’s produce a broader time profile.
The scintillator light yield and quenching are described with Birks’ law, which is tuned using the PMT hit spectra of the 214BiPo coincidence events, following the procedure described in the previous analysis [7]. A single light yield is tuned by matching MC simulations of 214Bi to data, while the Birks’ constant for ’s, ’s, and protons are each modeled with a distinct, tuned value. The newly-fit values are consistent with those of the previous analysis.
After these calibrations were performed, a residual nonuniformity of the energy scale was observed as a function of position nearer the AV. Empirical corrections for both data and simulation were created by fitting the mean energies of the selected 214BiPo in bins of the vertical position and the squared horizontal radius .
A bilinear interpolation function was used to provide a continuous correction throughout the volume of the AV. After applying the correction, the uncertainties on the energy scale and resolution were evaluated as differences between the data and simulation to be 1.8% and 4.4% (relative), respectively. These uncertainties were consistent between the ’s and ’s.
Antineutrino selection.
Nuclear reactors produce a large, pure flux of with energies up to around 10 MeV in the beta decays of nuclear fission products. These antineutrinos can be detected via the inverse beta decay (IBD) interaction on hydrogen atoms in the detector medium: . This process has a 1.8-MeV threshold and produces a positron, which quickly annihilates with an electron in the medium, depositing a total energy of in the detector. The neutron thermalizes and captures on hydrogen with a lifetime of about 200 s, producing a 2.2-MeV . The coincidence of these prompt and delayed events provides a distinct signal to identify reactor antineutrinos, greatly reducing backgrounds.
Reactor antineutrino candidates are selected using the criteria summarized in Table 1.
A likelihood ratio (LR) is used to provide additional suppression of accidental coincidences. Similar to the LR method used in the analysis of the SNO+ water phase [6], the probability density functions (PDFs) for reactor- IBDs come from MC simulations, and the PDFs for accidentals are constructed from a data sample of randomly-paired events that pass the prompt and delayed event criteria. The LR is then calculated from the products of the 2D PDF of versus and the 1D PDF of the delayed event . In addition, Bayesian priors are applied to account for the rates of reactor- IBDs and accidentals. A cut of -3.5 on the LR was determined by maximizing a ratio of signal (S) and background (B) counts: S/.
A suite of cuts designed to remove instrumental backgrounds is applied to all events. All data within the 20 seconds following an identified muon or high-energy event are removed to avoid cosmogenically-induced backgrounds. Additionally, any potential background from fast neutrons produced by muons in the external water is mitigated by removing events within 10 s of an event with 3 or more hit outward-looking PMTs. Event pairs that pass the 214BiPo selection criteria in Table 1 or have additional coincident events, are also removed.
After applying all selection criteria, 59 coincidence pairs are observed in the dataset.
The distributions of , , and delayed reconstructed energy of these coincidences are in good agreement with expectations for neutron captures, which were obtained by simulations of reactor- IBDs [11]. Figure 1 shows the position distribution of the selected pairs, which are found to be uniform throughout the detector, as expected.
Figure 1: Positions of selected prompt and delayed events.
Reactor antineutrinos.
Over 99% of the expected reactor antineutrino flux at SNO+ comes from reactors in North America, around 60% of which comes from Ontario’s three Canadian Deuterium Uranium (CANDU) reactor complexes at baselines of 240 km, 340 km, and 350 km.
Neutrino oscillations across these distances result in multiple dips in the detected prompt energy spectrum, which are well preserved given the large contribution of the CANDU reactors to the total flux.
The flux is modeled using the thermal power outputs of reactor cores as a function of time, the average fractions of the four dominant fissile isotopes, the average energy released per fission [12], and the emitted antineutrino energy spectra per isotope.
Monthly-averaged thermal powers from yearly IAEA reports [13, 14] are used for all reactors except the CANDU reactors for which we use the hourly electrical output provided by IESO’s Generator Output and Capability reports [15].
These two reports show an average difference of % in reactor power over a period of a year.
The fission fractions of 235U, 238U, 239Pu and 241Pu depend on reactor type and also vary over time, due to the depletion of isotopes and refueling cycles. Average values of (0.568, 0.078, 0.297, 0.057) [16] are used for the large number of pressurized water reactors (PWR) and boiling water reactors (BWR). The CANDU reactors use the pressurized heavy water reactor (PHWR) fission fractions (0.52, 0.05, 0.42, 0.01) [17], which are stabilized by continuous refueling and are averaged over multiple cores in each complex.
Finally, the antineutrino energy spectrum emitted for each isotope is given by the Huber-Mueller model [18, 19]. The flux of the isotope model is known to be biased and is corrected by scaling by a factor of 0.9450.007 to match the global average of reactor flux measurements [20].
Several uncertainties on the flux are carried over from the detailed studies made by the Daya Bay experiment [20]. Per core, they include a 2.4% uncertainty from the isotopic emission spectra, 0.6% from fission fraction uncertainties, 0.5% from power output, and several other subdominant contributions. Studies carried out by the SNO+ collaboration produced consistent numbers.
The uncertainties on the shape of the predicted energy spectrum [20] are expected to have a smaller impact than the detector-related uncertainties, and to be negligible.
The number of target protons available to produce IBDs is defined by the fiducial volume. An uncertainty on the size of the volume is determined from the position reconstruction and an uncertainty on the number of protons is inherited from the proton density in the liquid scintillator cocktail. The former is estimated to result in less than 1% uncertainty on the volume and the latter is computed from the scintillator mass density and the hydrogen mass fractions of all included compounds. The uncertainty on mass density from measurements and temperature variations was found to be less than 0.5%. Hydrogen mass fractions of the various carbon chains are very similar, and deviations many times larger than the manufacturer’s limits would be required to produce even a modest impact on the proton density. However, Daya Bay assigned a 0.92% uncertainty to their proton number, primarily based on deviations of these mass fractions from combustion measurements [20]. In the absence of a direct independent measurement for the SNO+ cocktail, we conservatively adopt the same uncertainty.
The coincidence selection efficiency is obtained from MC simulations of IBDs occurring in the full scintillator volume, and is found to be 70%. Within the fiducial radius of 5.7 m, the efficiency is 83%.
With subpercent biases resulting from position and energy reconstruction uncertainties, all aforementioned uncertainties total to produce a 3% uncertainty on the normalization of the reactor flux.
Neutrino oscillations.
More than 95% of the incoming flux is expected to travel entirely through the North American continental crust, which has a relatively constant density [21, 22]. Therefore, the electron antineutrino survival probability is calculated using a full three-flavor mixing framework, accounting for the effect of constant matter density [23]:
where and are the baseline and antineutrino energy, respectively. All other quantities are defined in Ref. [23] and depend on , , , , and the crust’s electron density (8.131023 cm-3).
The matter effect induces a change in the flux of or less, and similarly, an change is induced in the effective .
For and eV2 [24], approximately 100 reactor- IBD interactions are expected within the AV per year.
Geoneutrinos.
Antineutrinos up to about 3.3 MeV are produced by decays within the Earth. Only decays from the 238U and 232Th chains produce ’s with energy greater than the IBD threshold. Methods based on Ref. [25] are used to predict the expected geoneutrino flux at SNO+ in Terrestrial Neutrino Units (TNU), where 1 TNU corresponds to 1 IBD interaction per year per 1032 free protons. Assuming a constant survival probability of
and a radiogenic heat of 20 TW, our model predicts geo- IBD rates of 36.38.7 TNU and 9.72.3 TNU from the 238U and 232Th chains, respectively. This gives an expected rate of 27 geo- IBDs within the AV per year.
Recognizing uncertainty in the range of possible Earth radiogenic heat values and in the local geology, the total geo- IBD rate is fit without a direct constraint.
As in geoneutrino studies by KamLAND and Borexino, knowledge of Earth’s Th and U chondritic abundances [26] motivates applying a constraint on the ratio of U/Th geo- IBD event rates when fitting the data. In this analysis, we use the predicted fluxes from our model to derive a constraint on the U/Th ratio of 3.7 1.3, where the uncertainty comes from combining each of the U and Th flux prediction uncertainties.
() backgrounds.
The dominant background in this analysis is from () interactions with the natural 13C present in the scintillator. These interactions produce 16O and a neutron, which thermalizes and captures, mimicking the IBD delayed signal. The prompt event can proceed through three channels, each producing a distinct energy peak. The neutron can elastically scatter protons, producing a signal in the apparent energy range of roughly 0.5 MeV to 3.5 MeV, or it can inelastically scatter off a 12C, which emits a 4.4-MeV . Alternatively, the 16O can be produced in an excited state, de-exciting by producing either a 6-MeV or an electron-positron pair that deposits about 6 MeV.
The dominant source of decays in the detector is 210Po. The rate of these decays is measured over time with a fit of the ’s energy peak, which quenches to be around 0.4 MeV.
The () background from 210Po implanted on the surface of the AV is reduced to a negligible level by the fiducial volume selection.
The average rate within the fiducial volume over this dataset is 38 Hz, nearly a factor of 5 reduction in specific 210Po activity compared to the initial oscillation measurement, as described in Ref. [7].
The probability of the 210Po undergoing an () interaction is calculated by integrating the energy-dependent interaction cross section up to the 210Po energy, and then multiplying by the 13C number density, in the same way as Ref. [7]. This, combined with the average 210Po rate, gives an average expected () rate of 0.21 events/day.
Disagreements between total cross section measurements and particularly large uncertainties in branching ratios (ground state vs. excited state) warrant a conservative approach to () uncertainties [27]. The uncertainties from Ref. [7] are assigned to the normalizations of the three () interaction channels: for proton elastic scatters and the 12C inelastic scatter, 30% uncertainties are assigned, and for the excited 16O channel, a 100% uncertainty is assigned.
214BiPo-like background.
Immediately after scintillator filling operations, the data showed higher rates of radioactive backgrounds primarily due to ingress of radon, resulting in most of these backgrounds decaying away with the 222Rn half-life of 3.8 days.
During these high-background periods, an excess of coincidences was observed with delayed event energies just below the 2.2-MeV region of interest.
The and distributions of these coincidences were consistent with a correlated decay of time and distance similar to those of IBD and 214BiPo coincident events, and the position distribution was uniform.
The prompt energy distribution of these events is in good agreement with the 214Bi decay spectrum (Q = 3.3 MeV) that precedes a 214Po decay, which has an energy of 7.8 MeV, but quenches down to about 0.8 MeV in visible energy in the scintillator.
The half-life of the 214Po decay is 164 s, which is close to the approximately 210-s neutron capture time for IBDs.
Rare decays of 214Po are too low in energy to create the excess.
A likely explanation is that these tails arise from alpha-proton elastic scattering interactions, in which more scintillation light is produced than by the ’s alone due to the lesser quenching of protons. Data show that decays of 215Po exhibit a similar tail extending to higher energies.
Figure 2: Fit of extended delayed energy distribution. The model includes a data-driven kernel density estimate of the 214Po spectrum based on the observed spectrum of 215Po. 214Po backgrounds, which dominate the lower half of the spectrum, were increased in this plot relative to the analyzed data via the inclusion of data from the period immediately after scintillator fill operations, when radon levels in the detector were elevated.
For this analysis, a data-driven model of the 214Po energy spectrum was created using a kernel density estimation of the 215Po energy distribution.
Figure 2 shows that the model fits well to the delayed energy spectrum, which is dominated by 214Po in the lower half. The corresponding prompt energy is modeled as 214Bi.
The LR can help suppress this type of background due to the differences in the delayed energy distributions between 214Po and 2.2-MeV neutron capture ’s.
Across the dataset used in the present analysis, the model predicts
that 1.11.1 214BiPo coincidence events will pass the IBD selection criteria, which has a small effect on sensitivity estimates for [28].
Other backgrounds.
The rate of accidental coincidences is calculated using the measured rates of events passing the selection criteria for prompt and delayed event candidates. After applying all selection criteria, including the LR cut, the calculation gives an expectation of 0.3 coincidences in the entire 134.4-day dataset. Since it is based directly on data, the prediction has a negligible uncertainty.
Atmospheric neutrinos can undergo neutral- and charged-current interactions with the hydrogen and carbon present in the scintillator, which can produce neutrons and mimic the IBD signal.
A study of simulated atmospheric neutrino interactions at SNO+ was performed using the GENIE Monte Carlo generator [29] and found a negligible contribution relative to the expected IBD signal.
Cosmogenic muon interactions in the detector can produce neutrons and sources of () decays. These backgrounds are reduced to a negligible level by the muon veto cuts.
Spectral Analysis and Results.
The expected numbers of signal and background events are listed in Table 2 and the corresponding prompt energy spectra are shown in Fig. 3.
The total expected number of selected coincidences is 54.7, in reasonable agreement with the 59 coincidences observed in the data.
Two independent analyses were performed to cross-check this event selection and the following fits.
Table 2: Expected and fitted numbers of signal and background events, and the total number of observed events in 134.4 days of data. Expectations show only systematic uncertainties and the geo- IBD rate is unconstrained. The fits are with unconstrained or constrained oscillation parameters.
Expectation
Fit (Uncon.)
Fit (Con.)
Reactor- IBD
27.90.8
25.1
27.50.9
18.25.2
17.2
17.2
Geo- IBD
7.2
12.0
11.1
214BiPo-like
1.11.1
1.21.1
1.21.1
Accidental
0.30.0
0.30.0
0.30.0
Total
54.7
55.8
57.3
Observed
59
59
59
Figure 3: Energy distribution of prompt events and best-fit (unconstrained) predictions.
A binned extended likelihood fit was applied to the prompt energy distribution to find the best-fit values of and .
The data and fitted energy spectra are shown in Fig. 3.
Nuisance parameters were constrained with Gaussian penalty terms added to the likelihood function. These parameters include the normalizations on reactor- IBDs, () channels, and the geoneutrino U/Th ratio, as well as energy systematics.
The uncertainties on the prompt energy are dominated by reconstruction uncertainties. The energy systematics are modeled as uncertainties in the energy scale – both linear and non-linear – and in the energy resolution. Two independent factors parameterize the linear energy scale uncertainties of ’s/’s (1.8%) and protons (3%) to account for their different scintillation response. A nonlinear scaling following Birks’ law is applied as a variation in Birks’ constant (5.4%). Finally, energy resolution uncertainty is modeled as a Gaussian smearing with a standard deviation of 4.4% [28, 27].
A 2D visualization of the measurement significance is shown in Fig. 4 as a grid scan of the likelihood over and , profiling over all other parameters.
Performing a complete fit to account for correlations between all parameters, the best-fit values for the unconstrained oscillation parameters are = () 10-5 eV2 and = .
A nearly equally significant result for occurs at the mirror value across 0.5.
Combining the previous measurement from SNO+ [7] with the present result produces minimal change; therefore, it has not been included.
The result for agrees with the KamLAND result, (7.53 0.18) 10-5 eV2 [1], within 1.
It shows a slightly worse agreement of 1.3 with the combined solar result from Super-K, (6.10) 10-5 eV2 [5].
The fit is repeated assuming Gaussian constraints of eV2 [1] and [30].
The fitted number of geo- IBDs is 11.1, which corresponds to a geo- IBD rate of TNU.
The measured geoneutrino rate has a relatively large uncertainty, making it consistent with the range of existing Earth models allowed by the two previous measurements.
The uncertainty in the geo- IBD rate is dominated by a negative correlation with the proton scattering component of () and is not significantly improved by constraining the oscillation parameters.
The resulting best-fit oscillation parameters are = 10-5 eV2 and .
The tension between solar and reactor results for , after combining SNO+ and KamLAND, is slightly greater than 1.5.
Table 3 summarizes the results of the two different fits. All nuisance parameters are found to be in good agreement with expectations in both cases. Given that is largely determined by the flux and by the spectrum, the two parameters are minimally correlated, leaving little to gain by constraining one and fitting the other.
Figure 4: The likelihood of the prompt energy spectrum as a function of vs. with 1 and 2 contours drawn. The likelihood is also shown separately for each variable, fixing the other to the best-fit value. The slight asymmetry about 0.5 is expected and gives a second best-fit point for near the global result 0.307 [24].
Table 3: Best-fit values for oscillation parameters and geo- IBD rate. Results are reported with no constraints on oscillation parameters and , and with Gaussian constraints according to their current best measurements (see text).
Fit (Uncon.)
Fit (Con.)
Geo- IBD rate (TNU)
Future sensitivities.
The sensitivity of this measurement is still limited by statistics and SNO+ is continuing to take data. The detector has been loaded with a secondary fluor (bis-MSB), increasing the light collected by more than 50%, and loading of tellurium is scheduled to begin in 2025. New data have been obtained with the enhanced light collection and sensitivity will remain similar when filled with tellurium, allowing continued study of reactor antineutrino oscillations and geoneutrinos.
Development of a novel event classifier showing improved rejection of the () background, as discussed in Ref. [7], is ongoing.
Because the () proton scatter events have an energy range similar to geo- IBDs, reduction of the () background is predicted to significantly improve the sensitivity to the geo- IBD rate.
Summary.
The SNO+ collaboration reports the second-most precise measurement of , using reactor antineutrinos. We find = () 10-5 eV2
while also measuring = . When applying Gaussian constraints according to external values of and , we obtain a measurement of = and of the geo- IBD rate: TNU.
Acknowledgements.
Capital funds for SNO+ were provided by the Canada Foundation for Innovation and matching partners:
Ontario Ministry of Research, Innovation and Science,
Alberta Science and Research Investments Program,
Queen’s University at Kingston, and
the Federal Economic Development Agency for Northern Ontario.
This research was supported by
Canada:
the Natural Sciences and Engineering Research Council of Canada,
the Canadian Institute for Advanced Research,
the Ontario Early Researcher Awards,
the Arthur B. McDonald Canadian Astroparticle Physics Research Institute;
U.S.:
the Department of Energy (DOE) Office of Nuclear Physics,
the National Science Foundation and the DOE National Nuclear Security
Administration through the Nuclear Science and Security Consortium;
UK:
the Science and Technology Facilities Council and the Royal Society;
Portugal:
Fundação para a Ciência e a Tecnologia (FCT-Portugal);
Germany:
the Deutsche Forschungsgemeinschaft;
Mexico:
DGAPA-UNAM and Consejo Nacional de Ciencia y Tecnología;
China:
the Discipline Construction Fund of Shandong University.
We also thank SNOLAB and SNO+ technical
staff; the Digital Research Alliance of Canada; the
GridPP Collaboration and support from Rutherford
Appleton Laboratory; and the Savio computational cluster
at the University of California, Berkeley. Additional long-term
storage was provided by the Fermilab Scientific Computing
Division.
For the purposes of open access, the authors have applied a Creative Commons Attribution licence to any Author Accepted Manuscript version arising. Representations of the data relevant to the conclusions drawn here are provided within this paper and its supplemental material.